The following is a table of available special characters and how to call them. Please note that nothing is final on this. I would welcome suggestions for names, other characters needed . . ..

In order to call these characters, put the following command in your file.

                \documentstyle[phonetic]{article}

And in order to get this table do

                     latex phonetic-table.tex

\schwa e e  \hausab  b b  \hausaB  B  B
\thorn  p p  \hausad  D D  \hausaD  T  T
\glottal ? ?  \hausak  k k  \hausaK  K K
\esh  s s  \varg  G G  \planck  h̄
\yogh  x x  \udesc  u u  \roty  y  y
\eth  d d  \wedge  U U  \rotm  Σm Σm
\emgma M M  \varomega Φ Φ  \ibar  ī
\engma n n  \revD  Θ Θ  \vod  v̊
\enya  N N  \fj  F F  \ubar  u
\flap  f f  \openo  c c  \rotOmega Ω  Ω
\voicedh h h  \varopeno C C  \vari  i  i
\hookd J J  \vara  a a  \barj  j  j
\rotvara A A  \rotr  r r  \rotw  w
\epsi  ” ”  \barlambda Δ Δ  \riota  Γ  Γ
õ  \ut\{o\}  {o \atop ..} \td\{o\}  o \syl\{o\}
{⌢ \atop mn} \labvel\{mn\}  a b a\upbar b  Ψ c \rc\{c\}
o –  \hill{o}  o̊ \od{o}  m^{h} m\uplett{h}
– o  \ohill{o}  ´ ō \acbar{\′}{o}  ´– o \acarc{\′}{o}
z \upharpoon  — \uphk  – \downhk
˜ \dbox  ˝  \sbox
↑ — \rhookup  ↑ — \lhookup
↑ ; — \bigrhookup  ↑ ; — \biglhookup
↑ ; — \Bigrhookup  ↑ ; — \Biglhookup
↑ ; — \biggrhookup  ↑ ; — \bigglhookup

↑ ; — \Biggrhookup  ↑ ; — \Bigglhookup

and the greek (non math, nonitalic \palpha, \pbeta, …)

fffiflffi fflıȷ ` ´ ˇ ˘ ¯ ˚ ¸ oßæ œøÆŒØ ! greek (non math but italic {\palpha, \pbeta …})

fffiflffi fflıȷ ` ´ ˇ ˘ ¯ ˚ ¸ oßæ œøÆŒØ ! greek (math)

αβγδ ϵζη θικλ μνξoπρ στυϕχψω